# Evaluation for Omni MLLMs

### v2

@@ 2.6 Dataset Comparison
In Table 2, we compare OmniVideoBench with representative audio-video benchmarks. While AV-Odyssey (Gong et al., 2024) and OmniBench (Li et al., 2024c) operate on single images, OmniVideoBench targets substantially more challenging videos with durations ranging from a few seconds to 30 minutes. Recent benchmarks such as AVTrustBench (Chowdhury et al., 2025), DAVE (Radevski et al., 2025), and MMAU (Sakshi et al., 2024) begin emphasizing audio–video coordination, but typically focus on specific capabilities or short clips. AVHBench (Sung-Bin et al., 2024) also evaluates audiovisual consistency, yet its tasks remain centered on shorter videos and hallucination detection. OmniVideoBench, by contrast, expands the scope to diverse video types, broader temporal spans, and fine-grained cross-modal reasoning, capturing richer dependencies between audio and vision. Compared to Daily-Omni (Zhou et al., 2025) and WorldSense (Hong et al., 2025), which also utilize multi-domain videos, OmniVideoBench places greater emphasis on explicit audiovisual collaboration. For instance, disabling audio causes Gemini-2.0-Flash’s performance to plummet to random levels, indicating that visual-only cues are insufficient. Furthermore, Figure 4 shows that widely used models such as Qwen2.5-Omni-7B perform closer to random guessing on our benchmark, indicating that OmniVideoBench presents significantly greater challenges than existing multimodal datasets.
@@ 3.1 Baseline Models
Models Audio Type Video Duration Avg. Music Sound Speech (0,1] min (1,5] min (5,10] min (10,30] min Omni-Modal Language Models (With Visual and Audio) Gemini-3.0-Pro 52.81 55.17 64.13 62.42 66.18 57.02 59.76 61.79 Gemini-2.5-Pro 38.46 57.72 61.66 57.83 64.43 55.02 55.94 58.90 Gemini-3.0-Flash 49.45 50.34 56.69 58.43 55.10 55.90 52.29 55.10 Gemini-2.5-Flash 39.56 57.04 53.17 55.42 55.10 47.37 52.11 52.40 Gemini-2.0-Flash 29.67 40.27 43.21 49.40 43.15 41.05 34.87 41.50 Qwen3-Omni-30B-A3B 37.36 34.67 39.26 45.78 37.03 38.86 35.11 38.40 OmniVinci-9B 30.77 32.67 32.15 38.55 34.11 30.13 27.10 32.10 Baichuan-Omni-1.5 24.18 31.33 31.36 28.92 31.78 28.38 32.44 30.70 HumanOmni-7B 20.87 31.08 31.61 36.57 29.36 29.60 29.25 30.50 VITA-1.5-7B 25.27 28.57 31.49 31.33 27.41 30.57 33.97 30.50 MiniCPM-o 27.47 28.57 30.24 31.43 28.49 34.53 26.15 29.70 Qwen2.5-Omni-7B 23.07 25.33 30.70 41.57 27.41 25.33 26.72 29.30 VideoLLaMA2-7B 26.37 30.67 29.25 32.00 28.20 29.60 28.29 29.20 Omni-Modal Language Models (Visual Only) Gemini-2.0-Flash 25.27 36.67 30.99 33.73 35.86 32.75 22.48 31.30 Qwen2.5-Omni-7B 27.47 26.67 26.22 28.31 27.11 24.45 25.95 26.40 Visual Language Models (Visual Only) Qwen2.5-VL-32B 32.97 32.00 31.49 38.55 31.20 29.26 30.53 31.80 Qwen2.5-VL-7B 29.67 31.33 29.51 25.90 30.03 31.88 30.15 29.80 Qwen2.5-VL-72B 26.37 29.33 29.91 33.13 30.03 31.88 24.43 29.50 Baseline LLMs DeepSeek-V3.1 28.57 26.17 27.28 30.91 27.57 25.00 26.44 27.60
@@ 3.3 Further Analysis
Effect of Different Numbers of Frames.

### v3

@@ 2.6 Dataset Comparison
As shown in Table 2, we compare OmniVideoBench with representative audio-video benchmarks. While AV-Odyssey (Gong et al., 2024) and OmniBench (Li et al., 2024c) operate on single images, OmniVideoBench targets substantially more challenging videos with durations ranging from a few seconds to 30 minutes. Recent benchmarks (Chowdhury et al., 2025; Radevski et al., 2025; Sakshi et al., 2024; Sung-Bin et al., 2024) have begun to emphasize audio–video coordination; however, they typically focus on specific capabilities or short clips. Some also evaluate audiovisual consistency, yet their tasks remain primarily centered on shorter videos and hallucination detection. OmniVideoBench, by contrast, expands the scope to diverse video types, broader temporal spans, and fine-grained cross-modal reasoning, capturing richer dependencies between audio and vision. Some other audio-visual benchmarks (Zhou et al., 2025; Hong et al., 2025), do not fully achieve a natural integration of the two modalities, while OmniVideoBench addresses this limitation effectively. For instance, disabling audio causes Gemini-2.0-Flash’s performance to plummet to the near-random level, indicating that visual-only cues are insufficient. Furthermore, Figure 4 shows that widely used models such as Qwen2.5-Omni-7B perform closer to random guessing on our benchmark, indicating that OmniVideoBench presents significantly greater challenges than existing multimodal datasets.
@@ 3.1 Baseline Models
Models Audio Type Video Duration Avg. Music Sound Speech (0,1] min (1,5] min (5,10] min (10,30] min Omni-Modal Language Models (With Visual and Audio) Gemini-3.0-Pro 52.81 55.17 64.13 62.42 66.18 57.02 59.76 61.80 Gemini-2.5-Pro 38.46 57.72 61.66 57.83 64.43 55.02 55.94 58.90 Gemini-3.0-Flash 49.45 50.34 56.69 58.43 55.10 55.90 52.29 55.10 Gemini-2.5-Flash 39.56 57.04 53.17 55.42 55.10 47.37 52.11 52.40 Gemini-2.0-Flash 29.67 40.27 43.21 49.40 43.15 41.05 34.87 41.50 Qwen3-Omni-30B-A3B 37.36 34.67 39.26 45.78 37.03 38.86 35.11 38.40 OmniVinci-9B 30.77 32.67 32.15 38.55 34.11 30.13 27.10 32.10 Baichuan-Omni-1.5 24.18 31.33 31.36 28.92 31.78 28.38 32.44 30.70 HumanOmni-7B 20.87 31.08 31.61 36.57 29.36 29.60 29.25 30.50 VITA-1.5-7B 25.27 28.57 31.49 31.33 27.41 30.57 33.97 30.50 MiniCPM-o 27.47 28.57 30.24 31.43 28.49 34.53 26.15 29.70 Qwen2.5-Omni-7B 23.07 25.33 30.70 41.57 27.41 25.33 26.72 29.30 VideoLLaMA2-7B 26.37 30.67 29.25 32.00 28.20 29.60 28.29 29.20 Omni-Modal Language Models (Visual Only) Gemini-2.0-Flash 25.27 36.67 30.99 33.73 35.86 32.75 22.48 31.30 Qwen2.5-Omni-7B 27.47 26.67 26.22 28.31 27.11 24.45 25.95 26.40 Visual Language Models (Visual Only) Qwen2.5-VL-32B 32.97 32.00 31.49 38.55 31.20 29.26 30.53 31.80 Qwen2.5-VL-7B 29.67 31.33 29.51 25.90 30.03 31.88 30.15 29.80 Qwen2.5-VL-72B 26.37 29.33 29.91 33.13 30.03 31.88 24.43 29.50 Baseline LLMs DeepSeek-V3.1 28.57 26.17 27.28 30.91 27.57 25.00 26.44 27.60
@@
3.2 Human Performance
It is worth noting that we also report the human test performance as human-level boundaries. We invited 10 qualified annotators, including 8 graduate students experienced in multimodal research and 2 experts trained in music-related analysis. Before the evaluation, 50 questions were randomly selected for 10 testers to answer simultaneously. The fact that the difference in results was no more than 5 questions indicates that the manual evaluation has minimal deviation and is viable. We consolidated questions requiring musical knowledge and assigned them to two music experts. The remaining questions were divided to ensure roughly equal distribution of speech and sound audio samples per set, which were then evenly assigned to high-level personnel such as graduate students to obtain persuasive results. Manual responses had no time constraints, yielding a final accuracy rate of 82.69%. This demonstrates that existing models still fall significantly short of human capabilities.
@@ 3.4 Further Analysis
Open-ended QA vs. MCQ.